OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. Find w if P = 68 and l = 10. A. W = 25 B. W = 15 C. W = 28 D. W = 24

OpenStudy (farmdawgnation):

Have you done problems like this before?

OpenStudy (anonymous):

Yes, but i have been very sick and missed alot of school and need help

OpenStudy (farmdawgnation):

Alright, that's fine. So, they give you an equation with some variables and give you values for some of the variables, right? So have you tried plugging the values in and seeing what it looks like after that?

OpenStudy (anonymous):

When i look at this and read what you just said, i get confused, maybe you can draw it out so i can understand better?

OpenStudy (farmdawgnation):

Right, so they give you this equation... P = 2l + 2w Then they tell you P = 68 and l = 10. Right? So, then that means: 68 = 2(10) + 2w Do you understand why?

OpenStudy (anonymous):

Hold on let me read this for a second, And thank you for helping me.

OpenStudy (anonymous):

No i don't understand why.

OpenStudy (farmdawgnation):

Ok, so P, l and w are variables. They aren't *real* numbers that we can multiply, divide, add, and subtract with. The whole point of the problem is to figure out the *real* number that w is supposed to represent. To help us do that, your problem gave us real numbers for P and l.

OpenStudy (farmdawgnation):

So, then we can take the original equation and then plug in the values that they gave us for P and l.

OpenStudy (anonymous):

Oh ok.

OpenStudy (farmdawgnation):

Do you follow so far?

OpenStudy (anonymous):

Yes

OpenStudy (farmdawgnation):

Ok, so now we have. 68 = 2(10) + 2w What do you think we could do to make this easier to understand?

OpenStudy (anonymous):

2x10 = 20 so replace the 2(10) with 20?

OpenStudy (farmdawgnation):

YES! Very good!

OpenStudy (farmdawgnation):

So we have: 68 = 20 + 2w Now, what next?

OpenStudy (anonymous):

Im not sure...

OpenStudy (farmdawgnation):

Ok, that's fine. So now we need to try and figure out w. The goal is to get it by itself.

OpenStudy (farmdawgnation):

The first thing that's *easiest* to do is to get rid of any numbers not attached to the w. So, let's get rid of the 20 on the right side. Do you know how we could do that?

OpenStudy (anonymous):

Move it over?

OpenStudy (farmdawgnation):

Hm. Yes, but it's a little bit more than that right? We've got to actually *subtract* it from both sides of the equation.

OpenStudy (anonymous):

Ok. Well how are you going to take 20 from 2?

OpenStudy (anonymous):

Go negative?

OpenStudy (farmdawgnation):

You don't, when you're subtracting a term out this way it doesn't hit the 2.

OpenStudy (anonymous):

Oh ok.

OpenStudy (farmdawgnation):

Really, what you're doing is this: 68 - 20 = 20 + 2w - 20 Right?

OpenStudy (anonymous):

Ok, makes sense

OpenStudy (farmdawgnation):

Right, so we would say that the 20 and -20 on the right side are of the same power. Neither have a variable following them, so we combine them and the 20 disappears from that side. Since the 2 has a w attached to it, the 20 can't touch it.

OpenStudy (farmdawgnation):

Only multiplication and division can detach and attach variables to numbers.

OpenStudy (anonymous):

Ok

OpenStudy (farmdawgnation):

So, if we subtract 20 from both sides, what do we get?

OpenStudy (anonymous):

48= 20 + 2w?

OpenStudy (farmdawgnation):

Very close. The right side of that is wrong, however.

OpenStudy (farmdawgnation):

When we subtract 20, we do it to *both* sides. So that 20 shouldn't be there.

OpenStudy (anonymous):

Oh ok, so it's 48 + 2w?

OpenStudy (farmdawgnation):

Er, actually 48 = 2w

OpenStudy (anonymous):

Sorry that's what i meant i pressed shift

OpenStudy (farmdawgnation):

It's all good.

OpenStudy (farmdawgnation):

So then w're most of the way there right? We've just got to get rid of that 2 and we'll know what w means. Do you have an idea of what we should do next?

OpenStudy (anonymous):

48/2?

OpenStudy (farmdawgnation):

Yes. But, do remember that if you ever need to do this for an exam, you'll be expected to show the steps in the proper format. So you will need to write something like... \[\frac{48}{2}=\frac{2w}{2}\] \[24 = w\]

OpenStudy (farmdawgnation):

Good job. :)

OpenStudy (anonymous):

Ok Thank you so much.!

OpenStudy (farmdawgnation):

np. I've got to go now, but best of luck!